What's the slope intercept form of the equation of the line that is perpendicular to 9x+3y=4 and contains (-1,-5)
Answer by rfer(16322)

3y=9x+4
y=3x+4/3
----------------
y=-1/3x+b
-5=-1/3*-1+b
b=-6 1/3
------------
y=-1/3x-6 1/3